Solve x-y=1and x+2y=7

Answer by Cromlix(4381)

Hi there,
x - y = 1.......(1)
x + 2y = 7......(2)
Subtract (1) from (2)
-3y = -6
y = 2
Substitute y = 2 into
x - y = 1
x - 2 = 1
x = 1 + 2
x = 3
x = 3, y = 2.
Hope this helps :-)